March 2005

JMM

Matematik 5GE

Assignment 2 The only way to learn mathematics is to do mathematics

Problem 1 Exercise 3.3 p 25.

Problem 2 Exercise 5.7 p 82.

Problem 3 Problem 5-2 p 87

Problem 4 Problem 5-3 p 87

Problem 5 Problem 5-4 p 87

(Due: Tuesday April 5th )